Suppose Sarah opens a lemonade stand. She hires Aisha and Sadeem for $12 a week ($6 each). She “rents” pitchers and spoons from her friend mother for $5 a week and spend $20 a week on paper cups, lemons, sugar, and other materials. Before she opened the stand, she was earning $15 a week selling newspapers. She grosses $72 in revenues a week from lemonade sales. Calculate accounting costs, economic costs, accounting profit, and economic profit. Evaluate how Sarah’s lemonade business is doing.
1
Expert's answer
2016-12-20T11:35:12-0500
Accounting costs = 12+5+20=$37;
Economic costs = 12+5+20+15=$52;
Accounting profit = 72-37=$35;
Economic profit = 72-52=$20.
I think, Sarah’s business is profitable and cost effective, so it is necessary to continue because it also brings economic benefits for her.